# BalancEdit: Dynamically Balancing the Generality-Locality Trade-off in Multi-modal Model Editing

### v1

@@ 5 Limitations
Another significant limitation is the memory-augmented approach’s handling of multi-hop model editing. The key-based similarity search struggles to capture multi-hop queries that depend on newly introduced knowledge, often due to the ambiguity of real-world facts. For example, if the CEO of X (formerly Twitter) were to change to Elon Musk, it would be difficult to update the response to the question, ‘Which social app is headed by the leader of SpaceX?’ A potential solution to this problem could involve dynamically defining the fine-grained influence scope, which would allow for more precise adjustments to changes in real-world facts and their implications for multi-hop questions.
@@ 6 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix E Baselines
Appendix F Implementation Details
Training Specifications
@@ Appendix F Implementation Details
Hyperparameters
In our comparisons of Finetuning, MEND and GRACE, we explore learning rates of $1.0$, $1e^{-1}$, $1e^{-2}$, $1e^{-3}$, $1e^{-4}$, and $1e^{-5}$. We observe that Finetuning, Memory, and MEND perform best with $1e^{-2}$.
The choice of layer to edit is another hyperparameter for all editors. In all our editor comparisons, each editor modifies the same layer. For miniGPT-4, this is the dense layer of the llama block (llama_model.model.layers[31].mlp.up_proj), for BLIP2-OPT moder, it is the OPT decoder layer (opt_model.model.decoder.layers[31].fc2.weight). Recent work supporting the importance of selecting the correct layers to fine-tune corroborates this (Cheng et al., 2023). However, it’s important to note that the choice of layer is a practical hyperparameter: for comparison purposes, we ensure editors are compared when editing the same layers. For the distance function, we use the Euclidean distance if it is not explicitly mentioned.

### v2

@@ 5 Limitations
Another limitation is the memory-augmented approach’s handling of multi-hop model editing. The key-based similarity search struggles to capture multi-hop queries that depend on newly introduced knowledge, often due to the ambiguity of real-world facts. For example, if the CEO of X (formerly Twitter) were to change to Elon Musk, it would be difficult to update the response to the question, ‘Which social app is headed by the leader of SpaceX?’ A potential solution to this problem could involve dynamically defining the fine-grained influence scope, which would allow for more precise adjustments to changes in real-world facts and their implications for multi-hop questions.
@@ 6 Conclusion
Acknowledgments
The work is supported in part by the U.S. Office of Naval Research Award under Grant Number N00014-24-1-2668, and the National Science Foundation under Grants IIS-2316306 and CNS-2330215.
@@ Acknowledgments
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix E Baselines
Appendix F Training Specifications